OpenStudy (anonymous):

Golf scores for a country club member are: 72, 71, 70, 70, 70, 69, 69, 69, 69, 69, 68, 68, 68, 67, 67. Find the mean, median, and mode, in that order. A. 69.5, 69, 68 and 70 B. 69, 69, 70 C. 69.1, 69, 69 D. none of the above

OpenStudy (anonymous):

C. Mean = (sum of numbers)/(number of elements) = 1036/15 =69.1 Median = middle number = 69 Mode = most common number = 69